# Density control of multi-agent swarms via bio-inspired leader-follower plasticity

### v1

@@ 3.3 Stability analysis
In this section, we assess the stability properties of our control solution, exerted through the velocity field $u$ in (2) and the reactive term $q$ in (17). From Theorem 1, we know that, if $\rho^{L}>0$, global convergence of $\rho$ toward $\bar{\rho}$ is ensured. Theorem 4 instead gives conditions for $\rho^{L}$ to be striclty positive at steady-state. Hence, we now prove local stability of the solution whose existence is proved in Theorem 4. Let us recall the function $e=\bar{\rho}-\rho$, and define error functions
@@
Theorem 8 (Local stability).
Under the conditions of Theorem 4, error functions (10), (38a), and (38b) locally converge to 0 almost everywhere if
$\|\bar{\rho}_{x}(\cdot)\|_{2}<\frac{2D}{\|f(\cdot)\|_{2}}$ (39)
@@
$e_{t}(x,t)=-K\,e(x,t),\qquad\qquad$ (40a)
$e^{*}_{t}(x,t)=-a\,e^{*}(x,t)+b\,e(x,t)$ $-b\,e^{\eta}(x,t),$ (40b)
$e^{\eta}_{t}(x,t)+[e^{\eta}(x,t)(f*\bar{\rho})(x)]_{x}-$ $\left[e^{\eta}(x,t)(f*e)(x,t)\right]_{x}=$
$=De^{\eta}_{xx}(x,t)-[\bar{\eta}^{F}(x)$ $(f*e)(x,t)]_{x}.$ (40c)
The first two equations of the error system are linear and do not involve spatial derivatives. The third equation, however, is nonlinear and involves spatial derivatives.
By linearizing the last equation about the origin, we find
$e^{\eta}_{t}(x,t)+[e^{\eta}(x,t)(f*\bar{\rho})(x)]_{x}=$
$=De^{\eta}_{xx}(x,t)-[\bar{\eta}^{F}(x)$ $(f*e)(x,t)]_{x}.$ (41)
We substitute (13) into (3.3), yielding
$e^{\eta}_{t}(x,t)+[e^{\eta}(x,t)$ $(f*\bar{\rho})(x)]_{x}=De^{\eta}_{xx}(x,t)$
$+\mathrm{exp}(-Kt)\left[\bar{\eta}^{F}(x)v^{0}(x)\right]_{x},$ (42)
@@
We introduce the Lyapunov functional
$V(t)=\|e^{\eta}(\cdot,t)\|_{2}^{2}.$ (43)
The time derivative of $V$ can be expressed as
$V_{t}(t)=2\int_{\mathcal{S}}e^{\eta}(x,t)e^{\eta}_{t}(x,t)\,\mathrm{d}x=\\ =2D\int_{\mathcal{S}}e^{\eta}(x,t)e^{\eta}_{xx}(x,t)\,\mathrm{d}x\\ -2\int_{\mathcal{S}}e^{\eta}(x,t)\left[e^{\eta}(x,t)(f*\bar{\rho})(x)\right]_{x}\,\mathrm{d}x\\ -2\mathrm{exp}(-Kt)\int_{\mathcal{S}}e^{\eta}(x,t)\tilde{v}(x)\mathrm{d}x,$ (44)
where we substituted (3.3) and $\tilde{v}=(\bar{\eta}^{F}v^{0})_{x}$ We expand the fist term on the right-hand side of (44) as
$2D\int_{\mathcal{S}}e^{\eta}(x,t)e^{\eta}_{xx}(x,t)\,\mathrm{d}x=-2D\int_{\mathcal{S}}\left(e^{\eta}_{x}(x,t)\right)^{2}\,\mathrm{d}x=\\ =-2D\|e_{x}^{\eta}(\cdot,t)\|_{2}^{2},$ (45)
where we applied integration by parts (recalling the periodicity of the functions). We similarly expand the second term on the right-hand side of (44) as
$-2\int_{\mathcal{S}}e^{\eta}(x,t)\left[e^{\eta}(x,t)(f*\bar{\rho})(x)\right]_{x}\,\mathrm{d}x=\\ =2\int_{\mathcal{S}}e^{\eta}_{x}(x,t)e^{\eta}(x,t)(f*\bar{\rho})(x)\,\mathrm{d}x=\\ =\int_{\mathcal{S}}\left[\left(e^{\eta}(x,t)\right)^{2}\right]_{x}(f*\bar{\rho})(x)\,\mathrm{d}x=\\ =-\int_{\mathcal{S}}\left(e^{\eta}(x,t)\right)^{2}(f*\bar{\rho})_{x}(x)\,\mathrm{d}x,$ (46)
where we used integration by parts (twice), and exploited the identity $\left[\left(e^{\eta}\right)^{2}\right]_{x}=2e^{\eta}e^{\eta}_{x}$. Substituting (45) and (46) into (44), we obtain
$V_{t}(t)=-2D\|e_{x}^{\eta}$ $(\cdot,t)\|_{2}^{2}-\int_{\mathcal{S}}\left(e^{\eta}(x,t)\right)^{2}(f*\bar{\rho})_{x}(x)\,\mathrm{d}x$
$-2\mathrm{exp}(-Kt)\int_{\mathcal{S}}e^{\eta}(x,t)\tilde{v}(x)\mathrm{d}x.$ (47)
By using Poincaré-Wirtinger inequality (see Lemma 2 in [19]), we can bound this as
$V_{t}(t)\leq-2D\|e^{\eta}$ $(\cdot,t)\|_{2}^{2}-\int_{\mathcal{S}}\left(e^{\eta}(x,t)\right)^{2}(f*\bar{\rho})_{x}(x)\,\mathrm{d}x$
$-2\mathrm{exp}(-Kt)\int_{\mathcal{S}}e^{\eta}(x,t)\tilde{v}(x)\mathrm{d}x.$ (48)
For the second term on the right-hand side of (3.3), we have
$\left|\int_{\mathcal{S}}\left(e^{\eta}(x,t)\right)^{2}(f*\bar{\rho})_{x}(x)\,\mathrm{d}x\right|\leq\\ \leq\int_{\mathcal{S}}\left|\left(e^{\eta}(x,t)\right)^{2}(f*\bar{\rho})_{x}(x)\right|\,\mathrm{d}x=\\ =\|e^{\eta}(\cdot,t)e^{\eta}(\cdot,t)(f*\bar{\rho})_{x}(\cdot)\|_{1}\leq\\ \leq\|e^{\eta}(\cdot,t)\|_{2}\|e^{\eta}(\cdot,t)\|_{2}\|(f*\bar{\rho})_{x}(\cdot)\|_{\infty}\leq\\ \leq\|e^{\eta}(\cdot,t)\|_{2}^{2}\|f(\cdot)\|_{2}\|\bar{\rho}_{x}(\cdot)\|_{2},$ (49)
@@
$\left|-2\mathrm{exp}(-Kt)\int_{\mathcal{S}}e^{\eta}(x,t)\tilde{v}(x)\mathrm{d}x\right|\leq\\ \leq 2\mathrm{exp}(-Kt)\|e^{\eta}(\cdot,t)\tilde{v}(\cdot)\|_{1}\\ \leq 2\mathrm{exp}(-Kt)\|e^{\eta}(\cdot,t)\|_{2}\|\tilde{v}(\cdot)\|_{2}$ (50)
By applying bounds (49) and (50) to (3.3), we obtain
$V_{t}(t)\leq\left(-2D+\|f(\cdot)\|_{2}\|\bar{\rho}_{x}(\cdot)\|_{2}\right)V(t)\\ +2\mathrm{exp}(-Kt)\|\tilde{v}(\cdot)\|_{1}\sqrt{V(t)}.$ (51)
If $\|\bar{\rho}\|_{2}<2D/\|f\|_{2}$, the right-hand side in (51) converges to 0 thanks to Lemma 4 in [19] (with $\beta=-2D+\|f\|_{2}\|\bar{\rho}_{x}\|_{2}$ $\gamma=0$, and $\delta=2\|\tilde{v}\|_{2}$ ). Hence, by Comparison Lemma [42], we know $\|e^{\eta}\|_{2}^{2}$ (locally) converges to 0.
Since (40a) converges point-wise and (40c) converges locally in $\mathcal{L}^{2}(\mathcal{S})$, we can analyze (40b), rewriting it as
@@ 4 Numerical Validation
We validate our theoretical results through numerical simulations, demonstrating the effectiveness of our control strategy. We characterize performance through the time evolution of the Kullback-Leibler (KL) divergence between $\rho$ and $\bar{\rho}$, that is,
$\mathcal{D}_{KL}(t)=\int_{\mathcal{S}}\rho(x,t)\log\left(\frac{\rho(x,t)}{\bar{\rho}(x)}\right)\,\mathrm{d}x.$ (55)
As a steady-state performance index, we consider $\mathcal{D}_{KL}^{ss}=\mathcal{D}_{KL}(t_{\mathrm{f}})$, where $t_{\mathrm{f}}$ is the final instant of the simulation.
4.1 Bimodal regulation
Figure 2: Robustness analysis to perturbations in (a) diffusion coefficients and (b) parameters of the interaction kernels. For different values of $p$, we show $\mathcal{D}_{KL}^{ss}$ (blue) and leaders’ mass (orange) at steady-state (in solid gray the predicted minimum plasticity ensuring feasibility).
[…]
4.2 Robustness analysis
Next, we demonstrate how the fraction of plastic agents affects robustness with respect to perturbations. We consider a desired von Mises distribution ($\kappa=1$, $\mu=0$), Morse interactions as in (35)($L_{a}=\pi$, $L_{r}=\pi/4$, $\alpha=2$), $K=10$, and $D=0.02$. We introduce perturbations to either the diffusion coefficient – doubling it for the followers in Eqs. (1c) and (1b) with respect to its nominal value used for the leaders in (1a)– or the interaction kernel parameters – reducing $L_{a}$ by 20% and increasing $L_{r}$ by 20% with respect to their nominal values for all the agents. Starting from equilibrium configurations [^1] For values of $p$ below the minimum threshold prescribed by Theorem 4, steady-state configurations $\bar{\rho}^{L}$ and $\bar{\rho}^{F}$ are negative in some regions of the domain. For these cases, we translate initial configurations upwards to become non-negative and re-normalize them to a predefined mass., we assess performance degradation for different values of $p$. When the perturbation affects only followers (as for the test with respect to $D$), we choose $K_{LF}$ and $K_{FL}$ to ensure the steady-state leaders’ mass is constant across different values of $p$. This makes the amplitude of the perturbation constant when varying $p$. Results in Figs. 2(a) and 2(b) show that above the minimum threshold for $p$ predicted in Theorem 4, agents rearrange to counteract perturbations, maintaining steady-state performance. Below this threshold, performance degrades significantly.
@@ 5 Agent-based Model
To bridge the gap between our continuum theoretical framework and practical implementation, we develop a discrete agent-based model that captures the essential PDE dynamics at the individual agent level. This validates the continuum approximation and demonstrates implementability in realistic multi-agent systems.
The agent-based model consists of coupled stochastic differential equations describing agent positions, with stochasticity from Gaussian noise and role-switching according to reaction rates.
In particular, we choose
$\mathrm{d}x_{i}(t)=\Bigg{[}\frac{1}{N^{LF}}$ $\sum_{i=1}^{N^{LF}}f(\{x_{i}(t),x_{j}(t)\})$
$+\frac{1}{M}\sum_{i=1}^{M}$ $f(\{x_{i}(t),y_{j}(t)\})+u_{i}(t)\lambda_{i}(t)\Bigg{]}\mathrm{d}t$
$+\sqrt{2D}\,\mathrm{d}W_{i},\;i=1,\dots,N^{LF},$ (56a)
$\mathrm{d}y_{i}(t)=\Bigg{[}\frac{1}{M}\sum_{i=1}^{M}f($$\{y_{i}(t),y_{j}(t)\})$
$+\frac{1}{N^{LF}}$ $\sum_{i=1}^{N^{LF}}f(\{y_{i}(t),x_{j}(t)\})\Bigg{]}\mathrm{d}t$
$+\sqrt{2D}\,\mathrm{d}W_{i},\;i=1,\dots,M.$ (56b)
Here, $x_{i}$ are positions of leaders and plastic followers, $y_{i}$ are positions of non-plastic followers, $N^{LF}$ and $M$ are their respective numbers, and $W_{i}$ is a standard Wiener process. Control input $u_{i}(t)=u(x_{i},t)$ is computed via spatial sampling, and $\lambda_{i}\in\{0,1\}$ indicates if agent $i$ is a leader ($\lambda=1$) or follower ($\lambda=0$). The label $\lambda_{i}$ is updated stochastically: leaders switch to plastic followers with rate $\kappa_{LF}(x_{i},t)$, plastic followers switch to leaders with rate $\kappa_{FL}(x_{i},t)$, where such space- and time-dependent rates are found factorizing $q$ in (17) as
@@
Figure 3: Agent-based bimodal regulation. (a, b) Initial/final collective densities (solid black). In the inset, we show the discrete displacement of agents - leaders in blue, plastic followers in orange, non-plastic followers in purple (agents are plotted on concentric circles for visualization purposes). (c) Steady-state densities of the three populations (same color coding). (d) Time evolution of the KL divergence (top panel) and leaders’ and followers’ mass (bottom panel).
We consider a setup analogous to Sec. 4.1 with $N=1000$ agents (300 initial leaders, 300 initial plastic followers, 400 non-plastic followers), using Euler-Maruyama integration with $\Delta t=10^{-3}$ with $t_{\mathrm{f}}=15$. Control inputs $u$ and $q$ are computed via kernel density estimation from the agents’ positions (using the Matlab function circ_ksdensity with the msn method [43] and a width of 50 on a mesh of 600 points). Other parameters are set to $D=0.05$, $K=1$, $K_{FL}=2$, and $K_{LF}=1$.
In Fig. 3, we show a realization of (56). As one may expect, we register a performance degradation compared to the continuum case (see Fig. 1), due to the finite-size effect of the discretization. The steady-state leaders-to-followers ratio is approximately 1/2, consistent with Corollary 6. Due to the stochastic nature of (56), we repeat the same simulation scenario 50 times, recording an average KL divergence at steady-state of approximately 0.02 (±$\pm 0.01$), and an average leaders-to-followers mass ratio of 0.48 ($\pm$ 0.1).

### v2

@@ 3.3 Stability analysis
In this section, we assess the stability properties of our control solution, exerted through the velocity field $u$ in (11) and the reactive term $q$ in (15). From Theorem 1, we know that, if $\rho^{L}>0$, global convergence of $\rho$ toward $\bar{\rho}$ is ensured. Theorem 5 instead gives conditions for $\rho^{L}$ to be striclty positive at steady-state. Hence, we now prove global stability of the solution whose existence is proved in Theorem 5. Let us recall the function $e=\bar{\rho}-\rho$, and define error functions
@@
Theorem 10 (Global stability).
Under the conditions of Theorem 5, error functions (9), (37a), and (37b) globally converge to 0 a.e. if
$\|\bar{\rho}_{x}(\cdot)\|_{2}<{2D}/{\|f(\cdot)\|_{2}}$ (38)
@@
$e_{t}=-K\,e,$ (39a)
$e^{*}_{t}=-a\,e^{*}+b\,e-b\,e^{\eta},$ (39b)
$e^{\eta}_{t}+[e^{\eta}(f*\bar{\rho})]_{x}-[e^{\eta}(f*e)]_{x}=De^{\eta}_{xx}-[\bar{\eta}^{F}(f*e)]_{x}.$ (39c)
We substitute (12) into (39c), yielding
$e^{\eta}_{t}+[e^{\eta}(f*\bar{\rho})]_{x}{-\exp(-Kt)\left[e^{\eta}v^{0}\right]_{x}}\\ =De^{\eta}_{xx}{-}\exp(-Kt)\left[\bar{\eta}^{F}v^{0}\right]_{x},$ (40)
@@
We introduce the $\mathcal{L}^{2}$-norm of the error $V=\|e^{\eta}\|_{2}^{2}$, such that its time derivative can be expressed as
$V_{t}=2\int_{\mathcal{S}}e^{\eta}e^{\eta}_{t}\,dx=2D\int_{\mathcal{S}}e^{\eta}e^{\eta}_{xx}\,dx\\ {-2\exp(-Kt)\int_{\mathcal{S}}e^{\eta}\left[\tilde{v}-\left(e^{\eta}v^{0}\right)_{x}\right]dx}\\ -2\int_{\mathcal{S}}e^{\eta}\left[e^{\eta}(f*\bar{\rho})\right]_{x}\,dx,$ (41)
where we used (40) and $\tilde{v}=(\bar{\eta}^{F}v^{0})_{x}$ We expand the first term on the right-hand side of (41) as
$2D\int_{\mathcal{S}}e^{\eta}e^{\eta}_{xx}\,\mathrm{d}x=-2D\int_{\mathcal{S}}\left(e^{\eta}_{x}\right)^{2}\,\mathrm{d}x=-2D\|e_{x}^{\eta}\|_{2}^{2},$ (42)
where we applied integration by parts (recalling the periodicity of the functions). We similarly expand the last term on the right-hand side of (41) as
$2\int_{\mathcal{S}}e^{\eta}\left[e^{\eta}(f*\bar{\rho})\right]_{x}\,\mathrm{d}x=2\int_{\mathcal{S}}e^{\eta}_{x}e^{\eta}(f*\bar{\rho})(x)\,\mathrm{d}x\\ =\int_{\mathcal{S}}\left[\left(e^{\eta}\right)^{2}\right]_{x}(f*\bar{\rho})\,\mathrm{d}x=-\int_{\mathcal{S}}\left(e^{\eta}\right)^{2}(f*\bar{\rho})_{x}\,\mathrm{d}x,$ (43)
where we used integration by parts (twice), and exploited the identity $\left[\left(e^{\eta}\right)^{2}\right]_{x}=2e^{\eta}e^{\eta}_{x}$. We can apply the same procedure in (43) to establish
$2\exp(-Kt)\int_{\mathcal{S}}e^{\eta}\left(e^{\eta}v^{0}\right)_{x}\,\mathrm{d}x\\ =\exp(-Kt)\int_{\mathcal{S}}\left(e^{\eta}\right)^{2}v^{0}_{x}\,\mathrm{d}x.$ (44)
Substituting (42), (43) and (44) into (41), we obtain
$V_{t}=-2D\|e_{x}^{\eta}\|_{2}^{2}-\int_{\mathcal{S}}\left(e^{\eta}\right)^{2}(f*\bar{\rho})_{x}\,\mathrm{d}x\\ -2\mathrm{exp}(-Kt)\int_{\mathcal{S}}e^{\eta}\tilde{v}\,\mathrm{d}x{+\mathrm{exp}(-Kt)\int_{\mathcal{S}}(e^{\eta})^{2}{v}^{0}_{x}\mathrm{d}x}.$ (45)
By using Poincaré-Wirtinger inequality (see Lemma 2 in [7]), we can bound $V_{t}$ as
$V_{t}\leq-2D\|e^{\eta}\|_{2}^{2}-\int_{\mathcal{S}}\left(e^{\eta}\right)^{2}(f*\bar{\rho})_{x}\,\mathrm{d}x\\ -2\mathrm{exp}(-Kt)\int_{\mathcal{S}}e^{\eta}\tilde{v}\,\mathrm{d}x{+\mathrm{exp}(-Kt)\int_{\mathcal{S}}(e^{\eta})^{2}{v}^{0}_{x}\mathrm{d}x}.$ (46)
For the second term on the right-hand side of (46), we have
$\left|\int_{\mathcal{S}}\left(e^{\eta}\right)^{2}(f*\bar{\rho})_{x}\,\mathrm{d}x\right|\leq\int_{\mathcal{S}}\left|\left(e^{\eta}\right)^{2}(f*\bar{\rho})_{x}\right|\,\mathrm{d}x\\ =\|e^{\eta}e^{\eta}(f*\bar{\rho})_{x}\|_{1}\leq\|e^{\eta}\|_{2}\|e^{\eta}\|_{2}\|(f*\bar{\rho})_{x}\|_{\infty}\\ \leq\|e^{\eta}\|_{2}^{2}\|f\|_{2}\|\bar{\rho}_{x}\|_{2},$ (47)
@@
$\left|\int_{\mathcal{S}}\left(e^{\eta}\right)^{2}v^{0}_{x}\,\mathrm{d}x\right|\leq\|e^{\eta}e^{\eta}v^{0}_{x}\|_{1}\leq\|e^{\eta}\|_{2}^{2}\|v^{0}_{x}\|_{\infty}$ (48)
$\left|-2\mathrm{exp}(-Kt)\int_{\mathcal{S}}e^{\eta}\tilde{v}\,\mathrm{d}x\right|\leq 2\mathrm{exp}(-Kt)\|e^{\eta}\tilde{v}\|_{1}\\ \leq 2\mathrm{exp}(-Kt)\|e^{\eta}\|_{2}\|\tilde{v}\|_{2},$ (49)
Applying bounds (47), (48) and (49) to (46), yields
$V_{t}\leq\left(-2D+\|f\|_{2}\|\bar{\rho}_{x}\|_{2}\right)V{+\exp(-Kt)\|v^{0}_{x}\|_{\infty}V}\\ +2\mathrm{exp}(-Kt)\|\tilde{v}\|_{2}\sqrt{V}.$ (50)
If $\|\bar{\rho}\|_{2}<2D/\|f\|_{2}$, the right-hand side in (50) converges to 0 thanks to Lemma 4 in [7] (with $\beta=2D-\|f\|_{2}\|\bar{\rho}_{x}\|_{2}$ $\gamma={\|v^{0}_{x}\|_{\infty}}$, and $\delta=2\|\tilde{v}\|_{2}$). Hence, by the Comparison Lemma [35], $\|e^{\eta}\|_{2}^{2}$ converges to 0.
Since (39a) converges point-wise and (39c) converges globally in $\mathcal{L}^{2}(\mathcal{S})$, we can analyze (39b), rewriting it as
@@ 4 Numerical Validation
We validate our claims through two scenarios scoring performance via Kullback-Leibler (KL) divergence in time $\mathcal{D}_{KL}=\int_{S}\rho\log(\rho/\bar{\rho})\,\mathrm{d}x$. As a measure of steady-state performance, we use the KL divergence at the final instant of our simulations, that is $\mathcal{D}_{KL}^{\mathrm{ss}}=\mathcal{D}_{KL}(t_{\mathrm{f}})$. For the numerical integration of (1), we use central finite difference in space and forward Euler in time over a mesh of 600 grid points and a time step $\Delta t=10^{-3}$.
First, we consider a bimodal von-Mises distribution (two terms as in (37) with $\mu_{1}=\pi/2$, $\mu_{2}=-\pi/2$, $k=3$) with interactions driven by a Morse kernel (34) ($L_{a}=\pi$, $L_{r}=\pi/2$, $\alpha=2$), $D=0.05$, $K=1$, $K_{FL}=1$, $K_{LF}=2$, $\Phi^{F}=0.4$, $M^{L}(0)=M^{F}(0)=0.3$. Several messages can be gathered from the results in Fig. 1: ($i$) the proposed bio-inspired control scheme is successful in achieving a bimodal density distribution for the collective starting from a uniform one (see Figs. 1(a) and 1(b)), in agreement with Theorem 1 and 10; ($ii$) our choice of $\hat{p}$ based on Theorem 5, ensures the strict positivity of the steady-state density displacement of the leaders (see the inset in Fig.  1(b)); and (iiiiii) our choice of $K_{LF}$ and $K_{FL}$ ensures a steady-state leaders-to-followers mass ratio $\hat{r}=1/2$ in agreement with Corollary 7. We obtained similar results when deliberately violating the sufficient stability condition (38) pointing at the conservative nature of our result. These findings are omitted for brevity.
Next, we demonstrate how the fraction of plastic agents affects robustness. We consider a desired von Mises distribution ($\kappa=1$, $\mu=0$), Morse interactions as in (34)($L_{a}=\pi$, $L_{r}=\pi/4$, $\alpha=2$), $K=10$, and $D=0.02$. We introduce perturbations to either the diffusion coefficient – doubling it for the followers in Eqs. (1c) and (1b) with respect to its nominal value used for the leaders in (1a)– or the interaction kernel parameters – reducing $L_{a}$ by 20% and increasing $L_{r}$ by 20% with respect to their nominal values for all the agents. Starting from equilibrium configurations [^2] For $p<\hat{p}$ (see Theorem 5), steady-state configurations $\bar{\rho}^{L}$ and $\bar{\rho}^{F}$ are negative in some regions. For these cases, we translate initial configurations upwards to become non-negative and re-normalize them to a predefined mass., we assess performance degradation for different values of $p$. When the perturbation affects only followers (as for the test with respect to $D$), we choose $K_{LF}$ and $K_{FL}$ to ensure the steady-state leaders’ mass is constant across different values of $p$. This ensures that possible improved robustness properties are not due to an increase in the amount of leaders at steady-state. Results in Figs. 2(a) and 2(b) show that above the minimum threshold for $p$ predicted in Theorem 5, agents rearrange to counteract perturbations, maintaining steady-state performance. Below this threshold, performance degrades significantly.
Figure 2: Robustness analysis to perturbations in (a) diffusion coefficients and (b) parameters of the interaction kernels. For different values of $p$, we show $\mathcal{D}_{KL}^{ss}$ (blue) and leaders’ mass (orange) at steady-state (in solid gray the predicted minimum plasticity from Theorem 5).
@@ 5 Agent-based Model
To bridge the gap between our continuum framework and practical implementation, we develop a discrete agent-based model capturing essential PDE dynamics at the individual level. Following the approach in [20], where density dynamics of swarms with transient leaders are considered, the model consists of coupled stochastic differential equations with Gaussian noise and stochasticity due to roles’ switching
$\mathrm{d}x_{i}=\Bigg[\frac{1}{N^{LF}}\sum_{j=1}^{N^{LF}}f(\{x_{i},x_{j}\})+\frac{1}{M}\sum_{j=1}^{M}f(\{x_{i},y_{j}\})\bigg]\mathrm{d}t\\ +u_{i}\lambda_{i}\mathrm{d}t+\sqrt{2D}\,\mathrm{d}W_{i}{},\;i=1,\dots,N^{LF},$ (54a)
$\mathrm{d}y_{i}=\Bigg[\frac{1}{M}\sum_{j=1}^{M}f(\{y_{i},y_{j}\})+\frac{1}{N^{LF}}\sum_{j=1}^{N^{LF}}f(\{y_{i},x_{j}\})\Bigg]\mathrm{d}t\\ +\sqrt{2D}\,\mathrm{d}W_{i}{},\;i=1,\dots,M.$ (54b)
Here, $x_{i}$ are positions of leaders and plastic followers, $y_{i}$ are positions of non-plastic followers, $N^{LF}$ and $M$ are their respective numbers, and $W_{i}$ is a standard Wiener process. Control input $u_{i}(t)=u(x_{i},t)$ is computed via spatial sampling, and $\lambda_{i}\in\{0,1\}$ indicates if agent $i$ is a leader ($\lambda=1$) or follower ($\lambda=0$). $\lambda_{i}$ is updated stochastically with rates $\kappa_{LF}(x_{i},t)$ and $\kappa_{FL}(x_{i},t)$ – see (19).
@@
Figure 3: Agent-based bimodal regulation. (a, b) Initial/final collective densities (solid black). In the inset, discrete displacement of agents: leaders (blue), plastic followers (orange), non-plastic followers (purple), plotted on concentric circles for visualization. (c) Steady-state densities of the three populations. (d) Time evolution of KL divergence (top) and leaders’ and followers’ mass (bottom).
We consider a setup analogous to Sec. 4 with $N=1000$ agents (300 initial leaders, 300 plastic followers, 400 non-plastic followers), using Euler-Maruyama integration with $\Delta t=10^{-3}$, $t_{\mathrm{f}}=15$. Control inputs are computed via kernel density estimation using Matlab’s circ_ksdensity with msn method [36] (width 50, 600 mesh points). Parameters are set to $D=0.05$, $K=1$, $K_{FL}=2$, $K_{LF}=1$.
Fig. 3 shows a realization of (54). Performance degradation compared to the continuum case (Fig. 1) arises from finite-size effects. The steady-state leaders-to-followers ratio is approximately 1/2, consistent with Corollary 7. Over 50 stochastic realizations, we obtain average $\mathcal{D}_{KL}^{\mathrm{ss}}$ $0.02$ ($\pm 0.01$) and average leaders-to-followers mass ratio $0.48$ ($\pm 0.1$).